OpenStudy (donnie1999):

HELP PLEASE!!! Saidhari bought a a toaster oven for $90. This price was ¼ off the list price. What was the list price?

OpenStudy (donnie1999):

@aloud

OpenStudy (anonymous):

4x-x/4=90 3x/4=90 x=90*4/3 x=30*4 x=120 List price = $120